# How Can We Effectively Expand the Vocabulary of LLMs with 0.01GB of Target Language Text?

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable capabilities in many languages beyond English. Yet, LLMs require more inference steps when generating non-English text due to their reliance on English-centric tokenizers and vocabulary, resulting in higher usage costs to non-English speakers. Vocabulary expansion with target language tokens is a widely used cross-lingual vocabulary adaptation approach to remedy this issue. Despite its effectiveness in inference speedup, previous work on vocabulary expansion has focused on high-resource settings assuming access to a substantial amount of target language data to effectively initialize the embeddings of the new tokens and adapt the LLM to the target language. However, vocabulary expansion in low-resource settings has yet to be explored. In this paper, we investigate vocabulary expansion in low-resource settings by considering embedding initialization methods and continual pre-training strategies. Through extensive experiments across typologically diverse languages, tasks and models, we establish a set of strategies to perform vocabulary expansion for faster inference, maintaining competitive downstream performance to baselines with only 30K sentences ($\sim$0.01GB text data) from the target language. [^1] Our code and models are available on GitHub.
How Can We Effectively Expand the Vocabulary of LLMs
with 0.01GB of Target Language Text?
Atsuki Yamaguchi1, Aline Villavicencio1,2  and Nikolaos Aletras1 1School of Computer Science, University of Sheffield, United Kingdom 2Department of Computer Science, Institute of Data Science and Artificial Intelligence, University of Exeter, United Kingdom {ayamaguchi1,a.villavicencio,n.aletras}@sheffield.ac.uk
Figure 1: We explore various adaptation strategies for cross-lingual vocabulary expansion (Refer to caption) that allow inference speedups while retaining competitive performance to the source models in low-resource settings.
@@ 1 Introduction
Large language models (LLMs) have strong capabilities in English and other languages OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Groeneveld et al. (2024). Yet, processing non-English texts with LLMs is challenging. They suffer from tokenization overfragmentation [^2] See Appendix E.1 for a quantitative analysis. and thus require more inference steps due to the reliance on English-centric tokenizers and vocabulary, resulting in higher utility costs for non-English speakers Ahia et al. (2023); Petrov et al. (2023).
@@
Cross-lingual vocabulary adaptation (CVA) via vocabulary expansion has been proposed to adapt LLMs (including their tokenizers) to specific target languages. Vocabulary expansion approaches extend the vocabulary of a source model with tokens from a target language, followed by continual pre-training on target language data. A wide range of language-specific LLMs derived from an English-centric LLM such as Llama2 Touvron et al. (2023) have been made available following this approach, including Chinese Cui et al. (2023), Tamil Balachandran (2023), Portuguese Larcher et al. (2023), and Japanese Fujii et al. (2024) models, inter alia. Vocabulary expansion improves inference speed but often assumes access to a substantial amount of target language data for adaptation. For example, Chinese and Tamil LLaMAs make use of 20 and 12GB of target language text, respectively Cui et al. (2023); Balachandran (2023). Given the guaranteed large number of model updates in high-resource settings, the embeddings of new tokens are often simply randomly initialized, followed by the standard continual pre-training with a causal language modeling (clm) objective Cui et al. (2023); Balachandran (2023); Larcher et al. (2023); Choi et al. (2024). However, it is not clear how effective this approach to vocabulary expansion is in low-resource settings.
@@
In this paper, we seek to (i) answer if this widely used adaptation approach under high-resource settings is as effective in low-resource settings; and (ii) identify the best possible vocabulary expansion strategies for language adaptation in low-resource settings, maintaining similar performance to the source models with faster inference (Figure 1). Our key contributions are as follows:
@@
  • We present the first systematic study of vocabulary expansion-based adaptation of generative LLMs (i.e three English-centric models) in low-resource settings (i.e. assuming only 30K sentences, $\sim$0.01GB text data), on two generation tasks (i.e. machine translation and summarization) across ten typologically diverse languages.
@@
  • We find that vocabulary initialization approaches that use heuristics information from the source and target tokinizers are more effective. Furthermore, fine-tuning the top and bottom two layers of the LLM using a multi-token prediction objective Gloeckle et al. (2024) works better than fine-tuning the full model with clm. Finally, using a short input sequence length by splitting longer text into multiple sentences allows for a larger number of model updates, mitigating underfitting.
2 Related Work
Text Overfragmentation.
@@ 2 Related Work
LLMs tend to overfragment text in underrepresented languages Rust et al. (2021); Muller et al. (2021). Overfragmentation has significant implications for non-English speakers, including higher API costs Ahia et al. (2023); Petrov et al. (2023), slower inference Hofmann et al. (2022); Sun et al. (2023); Petrov et al. (2023), and lower downstream performance Bostrom and Durrett (2020); Rust et al. (2021); Toraman et al. (2023); Fujii et al. (2023).
Cross-lingual Vocabulary Adaptation.
@@
CVA methods have been proposed for adapting base LLMs to specific target languages for improving downstream performance and inference speed Cui et al. (2023); Balachandran (2023); Fujii et al. (2024); Yamaguchi et al. (2024). CVA via vocabulary expansion incorporates new tokens into the source vocabulary Chau et al. (2020); Cui et al. (2023); Balachandran (2023); Larcher et al. (2023); Fujii et al. (2024); Liu et al. (2024). CVA with vocabulary replacement replaces the entire source vocabulary with a new one from the target Minixhofer et al. (2022); Dobler and de Melo (2023); Ostendorff and Rehm (2023); Downey et al. (2023). More recent methods extend only the language modeling head Hong et al. (2024). CVA is followed by continual pre-training on target language data, often called language adaptive pre-training – LAPT Chau et al. (2020).
@@
Vocabulary expansion is widely used for developing language-specific generative LLMs from a source model, e.g. Chinese Cui et al. (2023) and Tamil Balachandran (2023) LLaMAs, and Swallow for Japanese Fujii et al. (2024). This line of work assumes access to a substantial amount of target language data, e.g. 312B characters of Japanese text used for Swallow. This might not be feasible in low-resource settings with limited target language data or compute. To the best of our knowledge, our work is the first to investigate vocabulary expansion for efficient LLM inference in extremely low-resource settings by assuming access to a small amount of target language data.
@@ 3 Problem Statement
Let $\mathcal{M}_{\text{s}}$ be a source LLM with a vocabulary $\mathcal{V}_{\text{s}}$. Given an additional target vocabulary $\mathcal{V}_{\text{new}}$, the aim is to learn a model $\mathcal{M}_{\text{t}}$ with the identical architecture as $\mathcal{M}_{\text{s}}$ that supports a target vocabulary $\mathcal{V}_{\text{t}}=\mathcal{V}_{\text{s}}\cup\mathcal{V}_{\text{new}}$. $\mathcal{V}_{\text{s}}$ and $\mathcal{V}_{\text{t}}$ have their corresponding tokenizer, $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$.
First, a target language-specific auxiliary tokenizer $\mathcal{T}_{\text{aux}}$ is trained on target language data $\mathcal{D}$. $\mathcal{T}_{\text{t}}$ is then derived from $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{aux}}$ along with their corresponding vocabularies $\mathcal{V}_{\text{s}}$ and $\mathcal{V}_{\text{aux}}$ by incorporating top $k$ frequent tokens ($\mathcal{V}_{\text{new}}$) that are in $\mathcal{V}_{\text{aux}}\backslash(\mathcal{V}_{\text{s}}\cap\mathcal{V}_{\text{aux}})$ into $\mathcal{V}_{\text{s}}$, resulting in $\mathcal{V}_{\text{t}}$. Subsequently, $\mathcal{M}_{\text{t}}$ is initialized with the weights of $\mathcal{M}_{\text{s}}$. Its embedding and output layer matrices are expanded such that the former is of dimensionality $|\mathcal{V}_{\text{t}}|\times H_{\text{t}}$ and the latter $H_{\text{t}}\times|\mathcal{V}_{\text{t}}|$, where $H_{\text{t}}$ is the hidden dimensionality of $\mathcal{M}_{\text{t}}$. Each new token representation is initialized by applying a target vocabulary initialization method (§4). If the weights of both matrices are not tied, they are initialized separately. Finally, $\mathcal{M}_{\text{t}}$ is adapted to the target language by training it on $\mathcal{D}$ using a clm objective.
4 Target Vocabulary Initialization
After vocabulary expansion, the new embeddings should be initialized. Our aim is to investigate the effectiveness and robustness of different initialization approaches in low-resource settings. For that purpose, we evaluate: (1) random initialization; (2) initialization based on auxiliary models; and (3) heuristic-based initialization.
@@ 4.1 Random Initialization
We randomly initialize new tokens by sampling from $\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ and $\sigma$ are the mean and standard deviation of token embeddings of $\mathcal{M}_{\text{s}}$ (Random). This is the most simple and common approach when adapting English-centric LLMs to a target language via vocabulary expansion in high-resource settings Cui et al. (2023); Balachandran (2023); Larcher et al. (2023); Choi et al. (2024).
@@ 4.2 Initialization Based on Auxiliary Models
FOCUS (Dobler and de Melo, 2023) is a state-of-the-art CVA method that relies on auxiliary embeddings, i.e. fastText Bojanowski et al. (2017), for initialization. [^3] Due to resource constraints, we only report results using FOCUS as a representative initialization method that relies on auxiliary embeddings since it outperforms other similar methods such as WECHSEL Minixhofer et al. (2022) for vocabulary replacement. The main assumption is that semantic transfer of embeddings should result in better initialization over Random. We apply FOCUS by tokenizing $\mathcal{D}$ using $\mathcal{T}_{\text{t}}$ and train a fastText model for each language. [^4] Preliminary experiments with FOCUS using off-the-shelf pre-trained word-level fastText models yielded lower performance (see Appendix E.2).
@@ 4.3 Heuristic-based Initialization
Mean.
A straightforward approach is to initialize each new token in $\mathcal{V}_{\text{new}}$ with the average of its corresponding source tokens obtained by using $\mathcal{T}_{\text{s}}$ Yao et al. (2021). Koto et al. (2021) and Gee et al. (2022) have also followed a similar approach for vocabulary replacement in domain adaptation.
Merge.
@@
Mean uses solely $\mathcal{T}_{\text{s}}$, which might produce subtokens from $\mathcal{V}_{\text{s}}$ that are not semantically related to the new target token. To overcome this issue, we propose using merge rules from $\mathcal{T}_{\text{t}}$ to effectively initialize $\mathcal{V}_{\text{new}}$. Merge rules describe how $\mathcal{T}_{\text{t}}$ can combine two subtokens into one, e.g. (‘a’, ‘t’) $\rightarrow$ ‘at’. Therefore, each new token in $\mathcal{V}_{\text{new}}$ can be decomposed into several existing subtokens from $\mathcal{V}_{\text{s}}$. We hypothesize that such subtokens can be more semantically related to the new token than those in Mean as they use information from $\mathcal{T}_{\text{t}}$.
@@
We first identify any merge rule in $\mathcal{T}_{t}$ whose resulting token is in $\mathcal{V}_{\text{new}}$. Subsequently, we generate a mapping from each new target token to source subtokens using the identified merge rules. Finally, for each new token, we compute the hierarchical mean of the associated source subtoken embeddings using the mapping information. Algorithm 1 in Appendix §A describes our method.
Align.
@@
The tokenization of a given word can change dynamically depending on the context, e.g. ‘the’ can be tokenized as (‘th’, ’e’) or (‘t’, ‘he’). Likewise, a new token in $\mathcal{V}_{\text{t}}$ is not always tokenized in the same way by $\mathcal{T}_{\text{s}}$ when operating on a sequence rather than a single token, which is the case for Mean. To obtain a more fine-grained semantic representation of a new token, we propose token alignment initialization that leverages mapping information between tokens tokenized with $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$ and the mapping frequency (i.e. more frequent mappings are more important for the final representation). This process allows us to consider different tokenization variants for the initialization of a given token.
For each sentence $x\in\mathcal{D}$, we first tokenize it using $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$. We then compare the two tokenized versions and generate a mapping list of tuples for each new token in $x$, e.g. a token $t_{5}$ in $\mathcal{V}_{\text{t}}$ is tokenized as [($t_{1}$, $t_{2}$), ($t_{3}$, $t_{4}$)] in $\mathcal{V}_{\text{s}}$. Subsequently, we concatenate all mapping lists of tuples in $\mathcal{D}$ for each new token $t\in\mathcal{V}_{\text{new}}$. For each $t$, we then construct a unique set of its constituent tuples, i.e. $\bm{T}_{t}=(\bm{t}_{1},\bm{t}_{2},\bm{t}_{3},\dots)$, and the corresponding frequency list $\bm{F}_{t}=(f_{\bm{t}_{1}},f_{\bm{t}_{2}},f_{\bm{t}_{3}},\dots)$, given the concatenated list in the previous step. Finally, we use $\bm{T}_{t}$ and $\bm{F}_{t}$ to aggregate the mapping information for token $t$ and generate its representation by computing $\sum_{\bm{t}\in\bm{T}_{t}}{\left[f_{\bm{t}}\frac{1}{|\bm{t}|}\sum_{t^{\prime}\in\bm{t}}{\bm{e}_{t^{\prime}}}\right]}$, where $\bm{e}_{t^{\prime}}$ is the vector representation (embedding or LM head) of the subtoken $t^{\prime}\in\mathcal{V}_{s}$.

### v3

@@ Abstract
Large language models (LLMs) have shown remarkable capabilities in many languages beyond English. Yet, LLMs require more inference steps when generating non-English text due to their reliance on English-centric tokenizers and vocabulary, resulting in higher usage costs to non-English speakers. Vocabulary expansion with target language tokens is a widely used cross-lingual vocabulary adaptation approach to remedy this issue. Despite its effectiveness in inference speedup, previous work on vocabulary expansion has focused on high-resource settings assuming access to a substantial amount of target language data to effectively initialize the embeddings of the new tokens and adapt the LLM to the target language. However, vocabulary expansion in low-resource settings has yet to be explored. In this article, we investigate vocabulary expansion in low-resource settings by considering embedding initialization methods and continual pre-training strategies. Through extensive experiments across typologically diverse languages, tasks and models, we establish a set of strategies to perform vocabulary expansion for faster inference, while striving to maintain competitive downstream performance to baselines. This is achieved with only 30K sentences ($\sim$0.01GB text data) from the target language. [^1] Our code and models are available via GitHub.
@@ 1 Introduction
Figure 1: We address the challenge of effectively expanding vocabulary for LLMs in low-resource settings. This is crucial for reducing inference steps when generating non-English text, as LLMs often rely on English-centric tokenizers and vocabulary. Our approach explores various adaptation strategies (Refer to caption) to achieve inference speedups while aiming to retain competitive performance. Our recommended strategy combines heuristic-based parameter initialization for new tokens with fine-tuning the top and bottom two layers of the model, using a short input sequence length and a multi-token prediction objective Gloeckle et al. (2024).
@@
Large language models (LLMs) have strong capabilities in English and other languages OpenAI (2023); OpenAI et al. (2024); Touvron et al. (2023); Jiang et al. (2023); Groeneveld et al. (2024); DeepSeek-AI et al. (2025). Yet, processing non-English texts with LLMs is challenging. They suffer from tokenization overfragmentation (see Figure 2 for quantitative analysis) and thus require more inference steps due to the reliance on English-centric tokenizers and vocabulary, resulting in higher utility costs for non-English speakers Ahia et al. (2023); Petrov et al. (2023); Ali et al. (2024).
@@
Cross-lingual vocabulary adaptation (CVA) via vocabulary expansion has been proposed to adapt LLMs (including their tokenizers) to specific target languages (Cui, Yang, and Yao, 2023; Fujii et al., 2024; Choi et al., 2024; Tejaswi, Gupta, and Choi, 2024; Mundra et al., 2024, inter alia). Vocabulary expansion approaches extend the vocabulary of a source model with tokens from a target language, followed by continual pre-training on target language data. A wide range of language-specific LLMs derived from an English-centric LLM such as Llama2 Touvron et al. (2023) have been made available following this approach, including Chinese Cui, Yang, and Yao (2023), Tamil Balachandran (2023), Portuguese Larcher et al. (2023), and Japanese Fujii et al. (2024) models, inter alia. Vocabulary expansion improves inference speed but often assumes access to a substantial amount of target language data for adaptation. For example, Chinese and Tamil Llamas make use of 20 and 12GB of target language text, respectively Cui, Yang, and Yao (2023); Balachandran (2023). Given the guaranteed large number of model updates in high-resource settings, the embeddings of new tokens are often randomly initialized, followed by the standard continual pre-training with a causal language modeling (clm) objective Cui, Yang, and Yao (2023); Balachandran (2023); Larcher et al. (2023); Choi et al. (2024). However, it is not clear how effective this approach to vocabulary expansion is in low-resource settings.
@@
In this article, we seek to (i) answer if this widely used adaptation approach under high-resource settings is as effective in low-resource settings; and (ii) identify the best possible vocabulary expansion strategies for language adaptation in low-resource settings, while striving to maintain similar performance to the source model with faster inference (Figure 1). Our key contributions are as follows:
@@
  • We present the first systematic study of vocabulary expansion-based adaptation of generative LLMs (i.e three English-centric models) in low-resource settings (i.e. assuming only 30K sentences, $\sim$0.01GB text data or up to approximately 5M tokens), on two generation tasks (i.e. machine translation and summarization) and two classification tasks (i.e. multiple-choice reading comprehension and general knowledge and reasoning) across ten typologically diverse languages.
@@
  • We find that target parameter initialization approaches that use heuristics information from the source and target tokenizers are more effective. Furthermore, fine-tuning the top and bottom two layers of the LLM using a multi-token prediction objective Gloeckle et al. (2024) works better than fine-tuning the full model with clm. Finally, using a short input sequence length by splitting longer text into multiple sentences allows for a larger number of model updates, mitigating underfitting.
  • To better understand holistic aspects of vocabulary expansion in low-resource settings, we conduct a range of analyses, including the extent of source (English) knowledge retention and a direct comparison with vocabulary replacement.
2 Background
2.1 Text Overfragmentation
@@
LLMs tend to overfragment text in underrepresented languages Rust et al. (2021); Muller et al. (2021). Overfragmentation has significant implications for non-English speakers, including higher API costs Ahia et al. (2023); Petrov et al. (2023), slower inference Hofmann, Schuetze, and Pierrehumbert (2022); Sun et al. (2023); Petrov et al. (2023), and lower downstream performance Bostrom and Durrett (2020); Rust et al. (2021); Toraman et al. (2023); Fujii et al. (2023); Ali et al. (2024).
To quantify this phenomenon, we calculate the average number of tokens on the FLORES-200 NLLB Team et al. (2022) dev set across our target languages and models (Figure 2). Following the established assumption that lower average token counts indicate more efficient tokenization Ahia et al. (2023), [^2] This assumption is valid specifically within the context of the FLORES-200 parallel machine translation corpus. our analysis reveals two key observations: (1) Burmese, Sinhala, and Telugu demonstrate the most severe overfragmentation among our ten target languages, requiring at least 6.8x more tokens than English across all models, and (2) English consistently exhibits the most efficient tokenization.
2.2 Cross-lingual Vocabulary Adaptation
@@
CVA methods have been proposed for adapting base LLMs to specific target languages for improving downstream performance and inference speed Cui, Yang, and Yao (2023); Balachandran (2023); Fujii et al. (2024); Yamaguchi, Villavicencio, and Aletras (2024). CVA via vocabulary expansion incorporates new tokens into the source vocabulary Balachandran (2023); Larcher et al. (2023); Pipatanakul et al. (2023); Lin et al. (2024); Cui, Yang, and Yao (2023); Kim, Choi, and Jeong (2024); Fujii et al. (2024); Choi et al. (2024); Nguyen et al. (2024); Tejaswi, Gupta, and Choi (2024); Mundra et al. (2024). CVA with vocabulary replacement replaces the entire or partial source vocabulary with a new one from the target Ostendorff and Rehm (2023); Csaki et al. (2023); Da Dalt et al. (2024); Remy et al. (2024); Yamaguchi, Villavicencio, and Aletras (2024); Dobler and de Melo (2024); Cahyawijaya et al. (2024). More recent methods include a hypernetwork for tokenizer transfer Minixhofer, Ponti, and Vulić (2024) and adapters for vocabulary alignment Han et al. (2025). CVA is typically followed by continual pre-training on target language data, often called language adaptive pre-training – LAPT Chau, Lin, and Smith (2020).
@@
Vocabulary expansion is widely used for developing language-specific generative LLMs from a source model, e.g. Chinese Cui, Yang, and Yao (2023) and Tamil Balachandran (2023) Llamas, and Swallow for Japanese Fujii et al. (2024). This line of work assumes access to a substantial amount of target language data, e.g. 312B characters of Japanese text used for Swallow. This might not be feasible in low-resource settings with limited target language data or compute. To the best of our knowledge, our work is the first to investigate vocabulary expansion for efficient decoder-based LLM inference in extremely low-resource settings by assuming access to a small amount of target language data (30K sentences).
@@ 3 Problem Statement
The goal in this article is to expand the vocabulary of a source LLM to effectively support a target language in low-resource settings. This involves transitioning from a source model $\mathcal{M}_{\text{s}}$ (with vocabulary $\mathcal{V}_{\text{s}}$ and tokenizer $\mathcal{T}_{\text{s}}$) to a model $\mathcal{M}_{\text{t}}$ that supports an expanded target vocabulary $\mathcal{V}_{\text{t}}$ and tokenizer $\mathcal{T}_{\text{t}}$.
More specifically, given $\mathcal{M}_{\text{s}}$, $\mathcal{V}_{\text{s}}$, $\mathcal{T}_{\text{s}}$, and target language data $\mathcal{D}$, $\mathcal{M}_{\text{t}}$ is constructed as follows:
  1. 1.
    Crafting the Auxiliary Tokenizer: A target language-specific auxiliary tokenizer is firstly trained on $\mathcal{D}$. This tokenizer is built from scratch using $\mathcal{D}$, allowing it to capture language-specific nuances. It comes with its own vocabulary, $\mathcal{V}_{\text{aux}}$.
  2. 2.
    Constructing the Target Vocabulary and Tokenizer:
    1. (a)
      The new tokens $\mathcal{V}_{\text{new}}$ are identified by taking the top $k$ most frequent tokens from $\mathcal{V}_{\text{aux}}$ that are not already present in $\mathcal{V}_{\text{s}}$ (i.e. $\mathcal{V}_{\text{new}}=\text{top }k\in\mathbb{N}\text{ tokens from }\mathcal{V}_{\text{aux}}\setminus(\mathcal{V}_{\text{s}}\cap\mathcal{V}_{\text{aux}})$).
    2. (b)
      The target vocabulary $\mathcal{V}_{\text{t}}$ is then formed by combining the original source vocabulary $\mathcal{V}_{\text{s}}$ with these newly identified tokens: $\mathcal{V}_{\text{t}}=\mathcal{V}_{\text{s}}\cup\mathcal{V}_{\text{new}}$.
    3. (c)
      The target tokenizer $\mathcal{T}_{\text{t}}$ is subsequently derived to operate on this expanded $\mathcal{V}_{\text{t}}$.
  3. 3.
    Initializing and Adapting the Target Model:
    1. (a)
      The target model $\mathcal{M}_{\text{t}}$ begins as a copy of $\mathcal{M}_{\text{s}}$, inheriting its identical architecture and pre-trained weights.
    2. (b)
      Its embedding and output layer matrices are then expanded to accommodate the larger $\mathcal{V}_{\text{t}}$. Specifically, the dimensionality of the embedding layer becomes $|\mathcal{V}_{\text{t}}|\times H_{\text{t}}$ and that of the output layer becomes $H_{\text{t}}\times|\mathcal{V}_{\text{t}}|$, where $H_{\text{t}}$ is the hidden dimensionality of $\mathcal{M}_{\text{t}}$.
    3. (c)
      Each representation for a new token (i.e. tokens in $\mathcal{V}_{\text{new}}$) is initialized using a target parameter initialization method (detailed in §4). If the weights of both embeddings and language modeling head (i.e. output layer) are not tied, they are initialized separately. [^3] We follow the original configuration of the source model regarding weight tying for the embeddings and output layer to preserve its original behavior as closely as possible.
    4. (d)
      $\mathcal{M}_{\text{t}}$ undergoes continual pre-training (i.e. LAPT) on the target language data $\mathcal{D}$ using a causal language modeling (clm) objective.
4 Target Parameter Initialization
After vocabulary expansion (step 2 in §3), the new embeddings should be initialized. Our aim is to investigate the effectiveness and robustness of different initialization approaches in low-resource settings. For that purpose, we evaluate: (1) random initialization; (2) initialization based on auxiliary models; and (3) heuristic-based initialization. [^4] Our primary focus in the subsequent discussion is on embedding initialization for simplicity. However, if the weights of the embeddings and the language modeling head are not tied, the language modeling head should also be initialized independently, using the identical process applied to the embeddings.
@@ 4.1 Random Initialization
For new tokens, we randomly initialize the weights of their embeddings by sampling from $\mathcal{N}(\mu,\sigma^{2})$. Here, $\mu$ and $\sigma$ are the mean and standard deviation of the token embeddings from $\mathcal{M}_{\text{s}}$ (Random). This is the most simple and common approach when adapting English-centric LLMs to a target language via vocabulary expansion in high-resource settings Cui, Yang, and Yao (2023); Balachandran (2023); Larcher et al. (2023); Choi et al. (2024).
@@ 4.2 Initialization Based on Auxiliary Models
FOCUS (Dobler and de Melo, 2023) is a state-of-the-art CVA method that relies on auxiliary embeddings, i.e. fastText Bojanowski et al. (2017), for initialization. [^5] Due to resource constraints, we only report results using FOCUS as a representative initialization method that relies on auxiliary embeddings since it outperforms other similar methods such as WECHSEL Minixhofer, Paischer, and Rekabsaz (2022) for vocabulary replacement. The main assumption is that semantic transfer of embeddings should result in better initialization over Random. We apply FOCUS by tokenizing $\mathcal{D}$ using $\mathcal{T}_{\text{t}}$ and train a fastText model for each language. [^6] Preliminary experiments with FOCUS using off-the-shelf pre-trained word-level fastText models yielded lower performance.
@@ 4.3 Heuristic-based Initialization
Mean
A straightforward approach is to initialize the weights of each new token in $\mathcal{V}_{\text{new}}$ by averaging the weights of their corresponding source tokens, which are identified using $\mathcal{T}_{\text{s}}$ Yao et al. (2021). Koto, Lau, and Baldwin (2021) and Gee et al. (2022) have also followed a similar approach for vocabulary replacement in domain adaptation. More recently, Tejaswi, Gupta, and Choi (2024) and Mundra et al. (2024) have demonstrated the effectiveness of this mean initialization for vocabulary expansion under high-resource settings (i.e. at least 200M tokens and 2.5B tokens, respectively). However, our study specifically investigates low-resource settings with only 30K sentences, equivalent to at most 5M tokens.
Merge
@@
Mean uses solely $\mathcal{T}_{\text{s}}$, which might produce subtokens from $\mathcal{V}_{\text{s}}$ that are not semantically related to the new target token. To overcome this issue, we propose using merge rules from $\mathcal{T}_{\text{t}}$ to effectively initialize $\mathcal{V}_{\text{new}}$. Merge rules describe how $\mathcal{T}_{\text{t}}$ can combine two subtokens into one. Through these rules, each new token in $\mathcal{V}_{\text{new}}$ can be decomposed into several existing subtokens from $\mathcal{V}_{\text{s}}$.
For instance, consider a new token: ‘superherohype’ in $\mathcal{V}_{\text{new}}$. According to the merge rules in $\mathcal{T}_{\text{t}}$, it can first be decomposed into (‘superhero’, ‘hype’). If ‘superhero‘ is also a new token, it can be further decomposed into (‘super’, ‘hero’). This process continues until all constituent parts are tokens from $\mathcal{V}_{\text{s}}$ (in this case, ’super’, ’hero’, and ’hype’). We hypothesize that such hierarchically derived subtokens from $\mathcal{V}_{\text{s}}$ are more semantically related to the new token than those obtained by simple averaging, as they leverage the specific tokenization information embedded in $\mathcal{T}_{\text{t}}$.
The initialization process is as follows:
  1. 1.
    We identify all merge rules in $\mathcal{T}_{\text{t}}$ that result in a token found in $\mathcal{V}_{\text{new}}$.
  2. 2.
@@
    Using these identified merge rules, we generate a hierarchical mapping for each new target token down to its constituent source subtokens in $\mathcal{V}_{\text{s}}$.
  3. 3.
    For each new token, we compute the hierarchical mean of the embeddings of its associated source subtokens, guided by the mapping information.
Align
Initializing the weights of new tokens in $\mathcal{V}_{\text{t}}$ by simply averaging the weights of their constituent subtokens from $\mathcal{T}_{\text{s}}$ as in Mean can be suboptimal. This naive approach does not account for how tokenization might change in a full sequence, as opposed to a single token.
Consider the word: ‘_cup’ and a new token: ’_cu’.777‘_’ stands for a whitespace. While Mean averages the weights of ‘_c’ and ‘u’ from $\mathcal{T}_{\text{s}}$ to initialize that of ’_cu’, this can be suboptimal. If $\mathcal{T}_{\text{s}}$ tokenizes ‘_cup’ as ‘_c’ + ‘up’, but $\mathcal{T}_{\text{t}}$ tokenizes it as ‘_cu’ + ‘p’, ‘_cu’ functions as a new, distinct subword unit. Its weight initialization should reflect this new role, perhaps focusing more on relevant overlapping components like ‘_c’.
@@
To obtain a more fine-grained semantic representation of a new token, we propose token alignment initialization that leverages mapping information between tokens tokenized with $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$ and the mapping frequency (i.e. more frequent mappings are more important for the final representation). This process allows us to consider different tokenization variants for the initialization of a given token.
Specifically, the weights of each new token $t\in\mathcal{V}_{\text{new}}$ are initialized as follows.
  1. 1.
    Generate Mappings from $\mathcal{D}$:
    1. (a)
      For each sentence $x\in\mathcal{D}$, we first tokenize it using both $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$.
    2. (b)
      We then compare these two tokenized versions to identify how each new token $t$ in $\mathcal{T}_{\text{t}}$’s output corresponds to sequences of subtokens from $\mathcal{T}_{\text{s}}$’s output. These correspondences are stored as a list of tuples, e.g. a token $t$ from $\mathcal{V}_{\text{t}}$ might map to $(t_{1},t_{2}),(t_{3},t_{4})]$ from $\mathcal{V}_{\text{s}}$ in different contexts.
  2. 2.
    Construct Unique Mappings and Frequencies:
    1. (a)
      We concatenate all such mapping lists from $\mathcal{D}$ for each new token $t$.
    2. (b)
      From this combined list, we construct a unique set of constituent tuples for $t$, denoted as $\mathbf{T}_{t}=(\mathbf{t}_{1},\mathbf{t}_{2},\mathbf{t}_{3},\dots)$, where each $\mathbf{t}_{i}$ is a sequence of subtokens from $\mathcal{V}_{\text{s}}$.
    3. (c)
      We also create a corresponding frequency list $\mathbf{F}_{t}=(f_{\mathbf{t}_{1}},f_{\mathbf{t}_{2}},f_{\mathbf{t}_{3}},\dots)$, indicating how often each unique tuple $\mathbf{t}_{i}$ appeared in the mappings for token $t$ on $\mathcal{D}$.
  3. 3.
    Compute Initial Representation: Finally, using $\mathbf{T}_{t}$ and $\mathbf{F}_{t}$, we aggregate the mapping information to generate the representation for token $t$ by computing $\sum_{\mathbf{t}\in\mathbf{T}_{t}}{\left[f_{\mathbf{t}}\frac{1}{|\mathbf{t}|}\sum_{t^{\prime}\in\mathbf{t}}{\mathbf{e}_{t^{\prime}}}\right]}$. Here, $\mathbf{e}_{t^{\prime}}$ is the embedding of the subtoken $t^{\prime}\in\mathcal{V}_{\text{s}}$, $f_{\mathbf{t}}$ is the frequency of mapping $\mathbf{t}$, and $|\mathbf{t}|$ is the number of subtokens in $\mathbf{t}$.